A source of sound frequency 550Hz emits waves of wavelength 600 mm in air at 20°C. What is the

velocity of sound in air at this temperature? What will be the wavelength of the sound from this source in air at 0°C?



1
Expert's answer
2021-03-04T13:17:25-0500

Given quantities:

"\lambda = 600mm =0.6m \space\space\space\space frequency \space\space f = 550Hz"

velocity of sound "v = f\lambda = 0.6*550 = 330 m/s"

We know, "v \backsim \sqrt{T}"

here, "T_1 = t_1 + 273 = 20+273 = 293K"

"T_2 = t_2 + 272 = 0+273 = 273K"

"\to \large\frac{vv_1}{v_2} = \sqrt{\frac{T_1}{T_2}} \to v_2 = v_1*\sqrt{\frac{T_2}{T_1}} =" "330*\large\frac{273}{293} =" "318.5m\/s"

"\lambda_2 = \large\frac{v_2}{f} = \frac{318.5}{550}" "= 0.579m\to 579mm"